how many ml of StarForce selective herbicide should be diluted into a 16 litre of water and at what area of land should solution covered?
1
Expert's answer
2019-09-30T10:49:22-0400

According to the website of StarForce manufacturers (see the link below), you need to use 700 ml of Star Force per one hectare. You need to add these 700 ml of herbicide to at least 10 liters of water. Write a proportion to figure out how many milliliters of the herbicide we need to dilute in 16 liters: if we need 700 ml for 10 l, we need "x" ml for 16 l, and the minimum required concentration provided by the manufacturers is 700 ml of herbicide in 10 liters of water, or 700/10=70 ml/l, i.e. 70 ml per 1 liter of water, must be the same in all volumes of water. In math it is called linear growth. Therefore, we can write a proportion


"\frac{700}{10}=\frac{x}{16},\\
\space\\
x=\frac{16}{10}\cdot700\text{ ml}=1120\text{ ml}."

How to solve a proportion? It's very simple if you apply a mnemonic rule:

to find the unknown, multiply the opposite numbers and divide this by the number opposite to unknown. Or multiply the diagonal, divide the vertical:


The same reasonings we can apply to the area: 700 ml are suitable for 1 ha according to the StarForce website, then 1120 ml is suitable for "y" hectares (again we face the linear growth, or the area grows by the same amount with each milliliter of the herbicide):


"\frac{700}{1}=\frac{1120}{y},\\
\space\\
x=\frac{1120}{700}\cdot1\text{ ha}=1.6\text{ ha}."





For more info about dosage see the StarForce bottle sticker: https://www.agro.basf.co.za/Documents/products_files/onkruiddoders_files/stellar__star_ENG.pdf